An aqueous solution of density d = 0.803 g / ml is prepared by dissolving 12.8 ml of propanol in 75 ml of water at 20 ° C. Calculate the percentage of propanol (a) by volume; (b) percentage by weight.


1
Expert's answer
2021-01-13T13:55:30-0500

density = 0.803g/ml

12.8ml of propanol in 75ml water

%v/v = "\frac{12.8}{75+12.8}" ×100


%v/v = "\frac{12.8}{87.8}" × 100 = 14.57 %


12.8ml propanol

mass of propanol = (12.8×0.803)g = 10.2784g

mass of water = 75g

%w/w = "\frac{10.2784}{75+10.2784}" ×100


%w/w = "\frac{10.2784}{85.2784}" × 100 = 12.05%